Find the Z score that correspond to 61.4% of the distribution area to the right
Answer #1

solution

Using standard normal table,

P(Z > z) = 61.4%

= 1 - P(Z < z) = 0.614  

= P(Z < z) = 1 - 0.614

= P(Z < z ) = 0.386

= P(Z < -0.29 ) = 0.0.386

z = -0.29 ( using z table )